# Aggregated demand flexibility prediction of residential thermostatically controlled loads and participation in electricity balance markets

### 22 Jan 2024 v2

@@ Abstract
The aggregate demand flexibility of a set of thermostatically controlled residential loads (TCLs) can be represented by a virtual battery (VB) in order to manage their participation in the electricity markets. For this purpose, it is necessary to know in advance and with a high level of reliability the maximum power that can be supplied by the aggregation of TCLs. A probability function of the power that can be supplied by a VB is introduced. This probability function is used to predict the demand flexibility using a new experimental probabilistic method based on a combination of Monte Carlo simulation and extremum search by bisection algorithm (MC&ESB). As a result, the maximum flexibility power that a VB can provide with a certain guaranteed probability is obtained. The performance and validity of the proposed method are demonstrated in three different case studies where a VB bids its aggregate power in the Spanish electricity balancing markets (SEBM).
keywords:
@@
Thermostatically controlled loads, Virtual batteries, Demand flexibility prediction, Monte Carlo simulation, Electricity balance markets
@@
[A1]organization=Centro Tecnológico CARTIF, addressline=Parque Tecnológico de Boecillo 205, postcode=47151, city=Boecillo, country=Spain
@@
[A2]organization=Instituto de las Tecnologías Avanzadas de la Producción, Universidad de Valladolid, addressline=Paseo Prado de la Magdalena s/n, city=Valladolid, postcode=47011, country=Spain
@@
[A3]organization=Departamento de Ingeniería Eléectrica, Universidad de Valladolid, addressline=Paseo Prado de la Magdalena s/n, city=Valladolid, postcode=47011, country=Spain
@@ A Bernouilli trial
1 Introduction
Electricity systems are undergoing a huge transformation due to the need to switch from fossil energy sources to variable renewable energy sources (RES). At the same time, rising energy prices are putting electricity markets as they are currently formulated to the test. The need to adapt the demand curve to mitigate price peaks requires managing demand flexibility. To achieve the desired transformation of the energy system, net metering and self-consumption must be encouraged, and regulation is beginning to change accordingly poullikkas2013review ; rehman2020penetration . At the same time, demand is allowed to participate in new power markets and others traditionally reserved for generation forouli2021assessment ; gellings2016demand .
In the residential sector, heating and cooling account for a large percentage of energy demand fleiter2016mapping ; sech2011analisis . Most of the appliances used for these purposes have a thermal inertia providing flexibility, that is, their electricity demand can be shifted over time without loss of utility. This is the case for thermostatically controlled loads (TCLs), such as refrigerators, electric water heaters and heat pumps. The aggregation of demand flexibility provided by many TCLs exhibits properties that resemble those of batteries, hence this aggregation is referred to as a virtual battery (VB). They have been extensively studied HHao1 ; SKhan ; martin2021flexibility .
VBs can be used to provide ancillary services zhao2016geometric and to participate in appropriate electricity markets song2019hierarchical . This requires accurate aggregate demand flexibility forecasting mechanisms. Several approaches have been proposed to forecast demand flexibility. In pinto2017multi an evolutionary particle swarm optimisation combined with support vector data description is used to find feasible trajectories for residential loads. In lucas2019load a combinatorial optimisation is proposed for flexibility prediction using non-intrusive load monitoring in the residential sector. Artificial neural networks are the main method employed in macdougall2016applying ; ponocko2018forecasting ; merce2020load . In hu2021flexibility , a temporal convolution network was applied to predict the flexibility of electric vehicle and domestic hot water systems. Furthermore, the use of regression models based on machine learning is discussed in ahmadiahangar2019residential . Finally, a comprehensive review of flexibility prediction methodologies is given in ahmadiahangar2020demand .
The main contribution of this work is the development of a novel method based on Monte Carlo simulation and extremum search using the bisection method (MC&ESB) to predict the flexibility that can be provided by a TCL aggregation modeled as a VB. In addition, a method to control the VBs and the variability of their behavior is studied. Finally, the Spanish electricity balancing market (SEBM) is used to test the potential of the proposed estimation method to facilitate demand flexibility aggregators to make offers in this type of markets.
The paper is organized as follows. Section 2 explains the TCLs and VBs models and their management methodology. Section 3 discusses the prediction of flexibility by introducing the supply power probability function of a VB and developing the MC&ESB method. The SEBM and the necessary requirements for demand participation are explained in Section 4. In Section 5, three case studies are presented; the first one estimates the likelihood function of a sample VB, the second one tests the MC&ESB method in three different scenarios in which VBs could participate in SEBM, and the third one evaluates the response of the VB controller to a requested power signal based on the SEBM results. Finally, Section 6 concludes the paper.
2 TCL and VB modelling and control
The aggregate flexibility of a set of TCLs can be represented by a VB. The management of the VB is enabled by the existence of a demand flexibility control system that manages the operation of each of the TCLs in the aggregation. In this work, the models and controller reported in martin2021flexibility are used due to the accuracy with which they are able to satisfy the power requirements of the system operator by measuring the current state of the appliances in real time. But, as explained below, some improvements have been implemented in the control system that allow better demand prediction and more flexibility in the aggregation response. The monitoring and communication system between TCLs and the controller is done using the architecture proposed in LAKSHMANAN2016705 .
2.1 The TCL model
A simple discrete-time model of a TCL is given by equation (1). It describes the time evolution of the internal temperature $\theta_{i}^{k}$, which depends on its binary status $u_{i}^{k}$ (ON, OFF), the forecast ambient temperature $\hat{\theta}_{a_{i}}^{k}$, and the disturbance $\omega_{i}^{k}$. The parameters of the model are: the thermal resistance $R_{th_{i}}$, the thermal capacity $C_{th_{i}}$, the nominal power $P_{i}$, and the performance coefficient $\eta_{i}$.
where
The average power that each TCL is expected to demand $P_{0_{i}}^{k}$ is calculated using equation (3).
The larger the safety band, the more flexibility the TCL will provide.
2.2 The VB model
Let $m_{+}^{k}$ (resp. $m_{-}^{k}$) denote the maximum charging power (resp. discharging power) that can be provided by the VB at each instant. These variables measure whether a given flexible power demand can be supplied by the TCL aggregation. They are given by equations (5) and (6), respectively.
$m_{+}^{k+1}$ $=\sum_{i=1}^{N}(1-2\phi_{i})\cdot(P_{i}-P_{0_{i}}^{k})-P_{+}^{k},$ (5)
$m_{-}^{k+1}$ $=\sum_{i=1}^{N}(1-2\phi_{i})\cdot P_{0_{i}}^{k}-P_{-}^{k},$ (6)
where $\phi_{i}$ is the device type, and $P_{+}^{k}$ and $P_{-}^{k}$ are the the sum of power of the TCLs which are unavailable for charging or discharging, respectively. See martin2021flexibility for a more detailed explanation of the VB model and its applications.
2.3 The demand flexibility control
Figure 1: VB controller diagram martin2021flexibility .
Additional improvements in the demand flexibility controller of martin2021flexibility have been included in this article. Firstly, the absolute error when tracking $r^{k}$ is reduced, as the power difference between switching on or off the last TCL selected in Priority-Based Dissagregation is checked. In addition, the number of $u_{i}^{k}$ changes are examined, and thus the activations of the short-cycle prevention constraint are reduced. Secondly, the controller allows to configure periods of time without demand management, which means that only Aggregation of TCLs step is executed. This is the case when no $r^{k}$ is required.
3 Demand flexibility prediction
3.1 The VB power supply probability function
Predicting the flexibility of VBs requires studying how the temperatures of TCLs are expected to change over time. However, this is not an easy task. The evolution of TCL temperatures is influenced by errors in ambient temperature forecasting, model inaccuracy and other disturbances related to the operation of the devices (e.g. when a refrigerator is opened). Representing these inaccuracies and perturbations requires the use of random models. The TCLs used in this work encapsulate all of this uncertainty in the perturbation parameter $\omega_{i}^{k}$. Usually, $\omega_{i}^{k}$ is considered to be distributed as a Gaussian random variable of zero expectation and constant variance $\sigma^{2}$, see mathieu2013energy .
$\Phi(x)=\mathbb{P}[S=1\mid x]$ (7)
Figure 2: MC&ESB method diagram martin2021flexibility .
Monte Carlo simulation is a probabilistic method useful for estimating a value under uncertainty, especially for complex systems papadopoulos2001uncertainty . It consists of evaluating a function a large number of times, so it finally converges to the most probable solution (the mean value) despite of the existing uncertainty zhang2021modern . The accuracy of the method increases as the number of evaluations of the function increases landau_binder_2014 .
$\mathbb{P}(\Xi=n\mid p)={N\binom{n}{m}n}p^{n}(1-p)^{N-n},\ p\in[0,1]$ (8)
$f(p\mid\Xi=n)$ $=\frac{\mathbb{P}(\Xi=n\mid p)f(p)}{\mathbb{P}(\Xi=n)}=\frac{{N\binom{n}{m}n}p^{n}(1-p)^{N-n}}{\mathbb{P}(\Xi=n)}$
$=\frac{1}{c(n,N)}p^{n}(1-p)^{N-n}$ (9)
[…]
$\hat{p}=\arg\max_{p\in[0,1]}f(p\mid\Xi=n)$ (10)
which is
$\hat{p}=\frac{n}{N}$ (11)
[…]
Theorem 1
Let $\epsilon$ and $\delta$ be scalars in the open unit interval $(0,1)$ then
$\mathbb{P}(p\geq 1-\epsilon\mid\Xi=N)\geq 1-\delta$ (13)
whenever
$N\geq\frac{\ln(1/\delta)}{\ln(1/(1-\epsilon))}-1$ (14)
Proof.
[…]
$=1-(1-\epsilon)^{N+1}$
In the Monte Carlo simulation stage of the MC&ESB method, the time evolution of the VB is simulated for each candidate value of power $x$. An estimate of the probability $p$ is given by Equation (11), where $N$ is obtained using Equation (14).
A schematic graphical description of the method is shown in Figure 2. The stage of the extremum search using the bisection algorithm in the MC&ESB method consists of searching for the maximum value of the power $x$ in absolute value that the VB can supply with probability 1. The method starts by defining a search interval, which must be between $[0,+\infty]$ when searching for flexibility in charging and between $[-\infty,0]$ when searching for flexibility in discharging. Next, the bisection algorithm computes a first value of $x$ and the estimate of the probability $p$ is computed by Equation (11). The bisection interval bounds $a$ and $b$ are updated at each iteration. The bisection algorithm terminates when $|a-b|<\gamma$, with $\gamma$ being a given tolerance. The solution found $x$ is the maximum flexible demand power that the VB can supply with probability greater than $1-\epsilon$ and confidence $1-\delta$.
4 Demand-side participation in SEBM
The MC&ESB method proposed in this paper provides demand aggregators with a way to determine how much power they could offer in electricity markets that allow demand-side participation. Several European countries have opened their markets to consumers participation, such us United Kingdom, Belgium, Germany or Sweden forouli2021assessment ; freire2022literature . Here, we focus on Spanish electricity balance markets (SEBM), managed by Red Eléctrica de España (REE). In these markets, active power bids can be either upward or downward (from the generation point of view). This is equivalent to a VB discharging, reducing consumption (negative flexibility) or charging, demanding more power (positive flexibility), respectively.
There are currently three SEBM: Secondary Regulation, Tertiary Regulation and Replacement Reserves BOEref .
All of them focus on maintaining the balance between generation and demand, and operate in fifteen-minute periods. The bids submitted to the markets consist mainly of the active power offered, the price, and some possible complex constraints, such as indivisibility. Once the bids are submitted, a matching algorithm is executed. The SEBM comply with the European Commission Regulation EUref and ENTSO-E nomenclature ENTSOEweb . The main characteristics of the three SEBM are given in Table 1.
Table 1: SEBM characteristics BOEref ; marsboom2018proposal .
ENTSO-E nomenclature
Automatic Frequency
Restoration Reserves (aFRR)
Manual Frequency
Restoration Reserves (mFRR)
Replacement Reserves (RR)
European platform PICASSO MARI TERRE
Mode of activation Automatic Manual Manual
Maximum ramping period Real-time 15 min 30 min
Minimum time in advance
for submission of bids
16:00 D-1 25 min 55 min
All the requirements that demand aggregators need to participate in SEBM are gathered in REEguia . According to this, the minimum power bid must be 1 MW.
5 Case Studies
This section presents three case studies. The first case estimates the power supply probability function describing the behavior of a VB composed of a set of TCLs. The second case tests the performance of the MC&ESB method for participating in SEBM in different scenarios. The third case shows how the control of a VB manages the aggregated set of TCLs to supply the power needed to bid in SEBM.
5.1 Case study 1: Estimation of VB supply probability function
The TCLs participating in the VB are 1000 refrigerators, 1000 electric water heaters and 1000 reversible heat pumps, which are considered as heating pumps in winter and cooling pumps in summer. Their characteristics are shown in Table 2. The location of the VB is Madrid, in Spain, and the season is summer. Thus, The forecast ambient temperature of every device $i$, $\hat{\theta}_{a_{i}}^{k}$, is considered constant and equal to $24$°C for refrigerators and electric water heaters, whereas for reversible heat pumps is the temperature of August, 10th at 15:00 UTC of the typical meteorological year, obtained with PVGIS huld2012new .
The accuracy and confidence parameters $\epsilon$ and $\delta$ have been set to 0.02 and 0.005, respectively. Then, applying Theorem 1, the number of trials is $N=262$. The MC&ESB method is applied to obtain $x_{\min}$ and $x_{\max}$, where $x_{\min}$ (resp. $x_{\max}$) is the minimum (resp. maximum) value of power that the VB can always supply for $N=262$ experiments. Therefore, the probability of supplying power $x$ for any $x\in[x_{\min},x_{\max}]$ is greater that $0.98$ with confidence at least of $0.995$. For this VB, $x_{\min}=-1138.3$ kW and $x_{\max}=5841.8$ kW. By analogy, the MC&ESB method is also applied to obtain $x^{\prime}_{\min}$ and $x^{\prime}_{\max}$, where $x^{\prime}_{\min}$ (resp. $x^{\prime}_{\max}$) is the limit value of power such that if $x\leq x^{\prime}_{\min}$ (resp. $x\geq x^{\prime}_{\max}$), the VB can never supply the power $x$ for $N=262$ experiments. Therefore, the probability of not supplying power $x$ for any $x\not\in[x^{\prime}_{\min},x^{\prime}_{\max}]$ is greater that $0.98$ with confidence at least of $0.995$. For this VB, $x^{\prime}_{\min}=-1296.2$ kW and $x^{\prime}_{\max}=5992.6$ kW. For $x\in[x^{\prime}_{\min},x_{\min}]$ (resp. $x\in[x_{\max},x^{\prime}_{\max}]$) the probability function $\Phi$ and their bounds for a confidence of at least $0.995$ are depicted in Figure 3. The curves have been obtained for $50$ equally spaced supply powers $x$. At each power $x$, the probability is estimated as $\hat{p}=n/N$, where $n$ is the number of experiments where the power $x$ is successfully supplied and $N=262$. The lower and upper bounds $\hat{p}-\epsilon_{1}$ and $\hat{p}+\epsilon_{2}$ have been obtained using equation (3.2) such that $\mathbb{P}[\hat{p}-\epsilon_{1}\leq p\leq\hat{p}+\epsilon_{2}]=1-\delta=0.995$. The bounds $\epsilon_{1}$ and $\epsilon_{2}$ have been selected to be equal, whenever possible. However, this is not possible when $\hat{p}$ approaches 00 or $1$. Consequently, non symmetric bounds are considered for these cases. The estimated demand probability function $\Phi$ and its confidence bounds can be used to characterize the capability of a VB. For example, for the VB of this study, we can state that it can supply a power $x\in[-1138.3,5841.8]$ during $15$ minutes with probability greater than $0.98$ and confidence $0.995$. Moreover, we can also state from the data represented in Figure 3 that the VB can supply a power $x=5900$ kW during $15$ minutes with probability $p\in[0.7021,0.8322]$ and confidence $0.995$. Thus, the knowledge of the power supply probability function $\Phi$ for a given VB is crucial to decide about participating in electrical markets.
Table 2: Range of values for the parameters of residential TCLs mathieu2012modeling (pump set-points from BOEref2 ).
Heating pumps $2$ $2.0$ $-5.6$ $3.5$ $22.0$ $0.5$
Cooling pumps $2$ $2.0$ $5.6$ $2.5$ $24.0$ $0.5$
Electric water heater $120$ $0.4$ $-4.5$ $1.0$ $48.5$ $3.0$
Refrigerator $90$ $0.6$ $0.3$ $2.0$ $2.5$ $1.5$
5.2 Case study 2: Demand flexibility prediction
Table 3: Flexibility power available at each scenario (kW).
Scenario 1 Scenario 2 Scenario 3
3608.0 -1482.9 -1250.1
The resulting flexibility power at each scenario is showed in Table 3. The highest amount of power flexibility is obtained in Scenario 1, Followed by Scenario 2 and Scenario 3. Usually, the VB can provide more positive than negative power flexibility because TCLs take longer to increase or decrease their temperature when this change is not caused by an electromechanical element martin2021flexibility . All calculated powers are above 1 MW in absolute value, which allows the aggregator to make bids in the SEBMs. These results demonstrate that the MC&ESB method is useful to calculate the availability of flexible power to participate in SEBMs and other similar markets in a short period of time, depending on the computational capabilities.
5.3 Case study 3: Demand flexibility control
Figure 5: Absolute error of the VB controller.
6 Conclusion
7 Funding
@@
8 CRediT authorship contribution statement
Alejandro Martín-Crespo: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Resources, Data Curation, Writing - Original Draft, Writing - Review & Editing, Visualization, Project administration. Enrique Baeyens: Conceptualization, Methodology, Validation, Formal analysis, Investigation, Writing - Review & Editing, Visualization, Supervision. Sergio Saludes-Rodil: Conceptualization, Methodology, Validation, Resources, Writing - Review & Editing, Supervision, Project administration, Funding acquisition. Fernando Frechoso-Escudero: Methodology, Writing - Review & Editing, Supervision.
9 Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
10 Data availability
Data will be made available on request.

### 05 Apr 2024 v3

@@ Abstract
The aggregate demand flexibility of a set of thermostatically controlled residential loads (TCLs) can be represented by a virtual battery (VB) in order to manage their participation in the electricity markets. For this purpose, it is necessary to know in advance and with a high level of reliability the maximum power that can be supplied by the aggregation of TCLs. A probability function of the power that can be supplied by a VB is introduced in this paper. This probability function is used to predict the demand flexibility using a rigorous experimental probabilistic method based on a combination of Monte Carlo simulation and extremum search by bisection algorithm. As a result, the maximum flexibility power that a VB can provide with a probabilistic guaranteed bound is obtained. The performance and validity of the proposed method are demonstrated and discussed in three different case studies where a VB bids its aggregate power in the Spanish electricity balancing markets (SEBM).
@@
Keywords: Thermostatically controlled loads; Virtual batteries; Demand flexibility prediction; Monte Carlo simulation; Electricity balance markets
1 Introduction
Electricity systems are undergoing a huge transformation due to the need to switch from fossil energy sources to variable renewable energy sources (RES). At the same time, rising energy prices are putting electricity markets as they are currently formulated to the test. The need to adapt the demand curve to mitigate price peaks requires managing demand flexibility. To achieve the desired transformation of the energy system, net metering and self-consumption must be encouraged, and regulation is beginning to change accordingly [1, 2]. At the same time, demand is allowed to participate in new power markets and others traditionally reserved for generation [3, 4].
In the residential sector, heating and cooling account for a large percentage of energy demand [5, 6]. Most of the appliances used for these purposes have a thermal inertia providing flexibility, that is, their electricity demand can be shifted over time without loss of utility. This is the case for thermostatically controlled loads (TCLs), such as refrigerators, electric water heaters and heat pumps. The aggregation of demand flexibility provided by many TCLs exhibits properties that resemble those of batteries, hence this aggregation is referred to as a virtual battery (VB). They have been extensively studied [7, 8, 9].
VBs can be used to provide ancillary services [10] and to participate in appropriate electricity markets [11]. This requires accurate aggregate demand flexibility forecasting mechanisms. Several approaches have been proposed to forecast demand flexibility. In [12] an evolutionary particle swarm optimisation combined with support vector data description is used to find feasible trajectories for residential loads. In [13] a combinatorial optimisation is proposed for flexibility prediction using non-intrusive load monitoring in the residential sector. Artificial neural networks are the main method employed in [14, 15, 16]. In [17], a temporal convolution network was applied to predict the flexibility of electric vehicle and domestic hot water systems. Furthermore, the use of regression models based on machine learning is discussed in [18]. Finally, a comprehensive review of flexibility prediction methodologies is given in [19].
The main contribution of this work is the development of a novel method based on Monte Carlo simulation and extremum search using the bisection method to predict the flexibility that can be provided by a TCL aggregation modeled as a VB. The key result for the development of the probabilistic method is formulated in Theorem 1 and allows estimating the probability of successfully supplying a given constant power demand response with a guaranteed confidence measure.
In addition, a method to control the VBs and the variability of their behavior is studied. Finally, the Spanish electricity balancing market (SEBM) is used to test the potential of the proposed estimation method to facilitate demand flexibility aggregators to make offers in this type of markets.
The paper is organized as follows. Section 2 explains the materials an methods, specifically the TCLs and VBs models and their management and control methodology. Besides, the electrical markets and the necessary requirements for demand participation are also explained. The main results are explained in detail in Section 3. Here, the flexibility prediction is characterized by introducing a probability function of the supply power of a VB. The probability of successfully supplying a given constant power during a demand response event with a measure of confidence is obtained by applying the result stated by Theorem 1. Based on this theoretical result, the MC&ESB method is developed, which can be easily applied by aggregators and energy service providers based on demand response to make energy bids in the electricity market with high confidence in their supply. In Section 4, three case studies are discussed; the first one estimates the likelihood function of a sample VB, the second one tests the MC&ESB method in three different scenarios in which VBs could participate in SEBM, and the third one evaluates the response of the VB controller to a requested power signal based on the SEBM results. Finally, Section 5 concludes the paper.
2 Materials and methods
The aggregate flexibility of a set of TCLs can be represented by a VB. The management of the VB is enabled by the existence of a demand flexibility control system that manages the operation of each of the TCLs in the aggregation. In this work, the models and controller reported in [9] are used due to the accuracy with which they are able to satisfy the power requirements of the system operator by measuring the current state of the appliances in real time. But, as explained below, some improvements have been implemented in the control system that allow better demand prediction and more flexibility in the aggregation response. The monitoring and communication system between TCLs and the controller is done using the architecture proposed in [20].
2.1 The TCL model
A simple discrete-time model of a TCL is given by Equation 1 and Equation 2. They describe the time evolution of the internal temperature $\theta_{i}^{k}$, which depends on its binary status $u_{i}^{k}$ (ON, OFF), the forecast ambient temperature $\hat{\theta}_{a_{i}}^{k}$, and the disturbance $\omega_{i}^{k}$. The parameters of the model are: the thermal resistance $R_{th_{i}}$, the thermal capacity $C_{th_{i}}$, the nominal power $P_{i}$, and the performance coefficient $\eta_{i}$.
where
The average power that each TCL is expected to demand $P_{0_{i}}^{k}$ is calculated using Equation 3.
The larger the safety band, the more flexibility the TCL will provide.
2.2 The VB model
Let $m_{+}^{k}$ (resp. $m_{-}^{k}$) denote the maximum charging power (resp. discharging power) that can be provided by the VB at each instant. These variables measure whether a given flexible power demand can be supplied by the TCL aggregation. They are given by Equation 5 and Equation 6, respectively.
$m_{+}^{k+1}$ $=\sum_{i=1}^{N}(1-2\phi_{i})\cdot(P_{i}-P_{0_{i}}^{k})-P_{+}^{k},$ (5)
$m_{-}^{k+1}$ $=\sum_{i=1}^{N}(1-2\phi_{i})\cdot P_{0_{i}}^{k}-P_{-}^{k},$ (6)
where $\phi_{i}$ is the device type, and $P_{+}^{k}$ and $P_{-}^{k}$ are the the sum of power of the TCLs which are unavailable for charging or discharging, respectively. See [9] for a more detailed explanation of the VB model and its applications.
2.3 The demand flexibility control
Figure 1: VB controller diagram [9].
Additional improvements in the demand flexibility controller of [9] have been included in this article. Firstly, the absolute error when tracking $r^{k}$ is reduced, as the power difference between switching on or off the last TCL selected in Priority-Based Dissagregation is checked. In addition, the number of $u_{i}^{k}$ changes are examined, and thus the activations of the short-cycle prevention constraint are reduced. Secondly, the controller allows to configure periods of time without demand management, which means that only Aggregation of TCLs step is executed. This is the case when no $r^{k}$ is required.
2.4 Demand-side participation in SEBM
Several European countries have opened their markets to consumers participation, such us United Kingdom, Belgium, Germany or Sweden [3, 21]. Here, we focus on Spanish electricity balance markets (SEBM), managed by Red Eléctrica de España (REE). In these markets, active power bids can be either upward or downward (from the generation point of view). This is equivalent to a VB discharging, reducing consumption (negative flexibility) or charging, demanding more power (positive flexibility), respectively.
There are currently three SEBM: Secondary Regulation, Tertiary Regulation and Replacement Reserves [22].
All of them focus on maintaining the balance between generation and demand, and operate in fifteen-minute periods. The bids submitted to the markets consist mainly of the active power offered, the price, and some possible complex constraints, such as indivisibility. Once the bids are submitted, a matching algorithm is executed. The SEBM comply with the European Commission Regulation [23] and ENTSO-E nomenclature [24]. The main characteristics of the three SEBM are given in Table 1.
Table 1: SEBM characteristics [22, 25].
ENTSO-E nomenclature
Automatic Frequency
Restoration Reserves (aFRR)
Manual Frequency
Restoration Reserves (mFRR)
Replacement Reserves (RR)
European platform PICASSO MARI TERRE
Mode of activation Automatic Manual Manual
Maximum ramping period Real-time 15 min 30 min
Minimum time in advance
for submission of bids
16:00 D-1 25 min 55 min
All the requirements that demand aggregators need to participate in SEBM are gathered in [26]. According to this, the minimum power bid must be 1 MW.
The method developed in this paper encourages demand-side participation because it provides demand aggregators with a rigourous tool to calculate how much power they could offer in electricity markets.
3 Results
3.1 The VB power supply probability function
Predicting the flexibility of VBs requires studying how the temperatures of TCLs are expected to change over time. However, this is not an easy task. The evolution of TCL temperatures is influenced by errors in ambient temperature forecasting, model inaccuracy and other disturbances related to the operation of the devices (e.g. when a refrigerator is opened). Representing these inaccuracies and perturbations requires the use of random models. The TCLs used in this work encapsulate all of this uncertainty in the perturbation parameter $\omega_{i}^{k}$. Usually, $\omega_{i}^{k}$ is considered to be distributed as a Gaussian random variable of zero expectation and constant variance $\sigma^{2}$, see [27].
$\Phi(x)=\mathbb{P}[S=1\mid x]$ (7)
Figure 2: MC&ESB method diagram [9].
Monte Carlo simulation is a probabilistic method useful for estimating a value under uncertainty, especially for complex systems [28]. It consists of evaluating a function a large number of times, so it finally converges to the most probable solution (the mean value) despite of the existing uncertainty [29]. The accuracy of the method increases as the number of evaluations of the function increases [30].
$\mathbb{P}(\Xi=n\mid p)={N\binom{n}{m}n}p^{n}(1-p)^{N-n},\ p\in[0,1]$ (8)
$f(p\mid\Xi=n)$ $=\frac{\mathbb{P}(\Xi=n\mid p)f(p)}{\mathbb{P}(\Xi=n)}$
$=\frac{{N\binom{n}{m}n}p^{n}(1-p)^{N-n}}{\mathbb{P}(\Xi=n)}$
$=\frac{1}{c(n,N)}p^{n}(1-p)^{N-n}$ (9)
@@ 3.2 The MC&ESB method for estimating a VB power supply probability function
$\hat{p}=\arg\max_{p\in[0,1]}f(p\mid\Xi=n)$ (10)
which is
$\hat{p}=\frac{n}{N}$ (11)
@@
Theorem 1
Let $\epsilon$ and $\delta$ be scalars in the open unit interval $(0,1)$ then
$\mathbb{P}(p\geq 1-\epsilon\mid\Xi=N)\geq 1-\delta$ (13)
whenever
$N\geq\frac{\ln(1/\delta)}{\ln(1/(1-\epsilon))}-1$ (14)
Proof.
[…]
$=(N+1)\left[\frac{p^{N+1}}{N+1}\right]_{1-\epsilon}^{1}$
$=1-(1-\epsilon)^{N+1}$
In the Monte Carlo simulation stage of the MC&ESB method, the time evolution of the VB is simulated for each candidate value of power $x$. An estimate of the probability $p$ is given by Equation 11, where $N$ is obtained using Equation 14.
A schematic graphical description of the method is shown in Figure 2. The stage of the extremum search using the bisection algorithm in the MC&ESB method consists of searching for the maximum value of the power $x$ in absolute value that the VB can supply with probability 1. The method starts by defining a search interval, which must be between $[0,+\infty]$ when searching for flexibility in charging and between $[-\infty,0]$ when searching for flexibility in discharging. Next, the bisection algorithm computes a first value of $x$ and the estimate of the probability $p$ is computed by Equation 11. The bisection interval bounds $a$ and $b$ are updated at each iteration. The bisection algorithm terminates when $|a-b|<\gamma$, with $\gamma$ being a given tolerance. The solution found $x$ is the maximum flexible demand power that the VB can supply with probability greater than $1-\epsilon$ and confidence $1-\delta$.
4 Discussion
This section discusses three case studies. The first case estimates the power supply probability function describing the behavior of a VB composed of a set of TCLs. The second case tests the performance of the MC&ESB method for participating in SEBM in different scenarios. The third case shows how the control of a VB manages the aggregated set of TCLs to supply the power needed to bid in SEBM.
4.1 Case study 1: Estimation of VB supply probability function
The TCLs participating in the VB are 1000 refrigerators, 1000 electric water heaters and 1000 reversible heat pumps, which are considered as heating pumps in winter and cooling pumps in summer. Their characteristics are shown in Table 2. The location of the VB is Madrid, in Spain, and the season is summer. Thus, The forecast ambient temperature of every device $i$, $\hat{\theta}_{a_{i}}^{k}$, is considered constant and equal to $24$°C for refrigerators and electric water heaters, whereas for reversible heat pumps is the temperature of August, 10th at 15:00 UTC of the typical meteorological year, obtained with PVGIS [31].
The accuracy and confidence parameters $\epsilon$ and $\delta$ have been set to 0.02 and 0.005, respectively. Then, applying Theorem 1, the number of trials is $N=262$. The MC&ESB method is applied to obtain $x_{\min}$ and $x_{\max}$, where $x_{\min}$ (resp. $x_{\max}$) is the minimum (resp. maximum) value of power that the VB can always supply for $N=262$ experiments. Therefore, the probability of supplying power $x$ for any $x\in[x_{\min},x_{\max}]$ is greater that $0.98$ with confidence at least of $0.995$. For this VB, $x_{\min}=-1138.3$ kW and $x_{\max}=5841.8$ kW. By analogy, the MC&ESB method is also applied to obtain $x^{\prime}_{\min}$ and $x^{\prime}_{\max}$, where $x^{\prime}_{\min}$ (resp. $x^{\prime}_{\max}$) is the limit value of power such that if $x\leq x^{\prime}_{\min}$ (resp. $x\geq x^{\prime}_{\max}$), the VB can never supply the power $x$ for $N=262$ experiments. Therefore, the probability of not supplying power $x$ for any $x\not\in[x^{\prime}_{\min},x^{\prime}_{\max}]$ is greater that $0.98$ with confidence at least of $0.995$. For this VB, $x^{\prime}_{\min}=-1296.2$ kW and $x^{\prime}_{\max}=5992.6$ kW. For $x\in[x^{\prime}_{\min},x_{\min}]$ (resp. $x\in[x_{\max},x^{\prime}_{\max}]$) the probability function $\Phi$ and their bounds for a confidence of at least $0.995$ are depicted in Figure 3. The curves have been obtained for $50$ equally spaced supply powers $x$. At each power $x$, the probability is estimated as $\hat{p}=n/N$, where $n$ is the number of experiments where the power $x$ is successfully supplied and $N=262$. The lower and upper bounds $\hat{p}-\epsilon_{1}$ and $\hat{p}+\epsilon_{2}$ have been obtained using Equation 3.2 such that $\mathbb{P}[\hat{p}-\epsilon_{1}\leq p\leq\hat{p}+\epsilon_{2}]=1-\delta=0.995$. The bounds $\epsilon_{1}$ and $\epsilon_{2}$ have been selected to be equal, whenever possible. However, this is not possible when $\hat{p}$ approaches 00 or $1$. Consequently, non symmetric bounds are considered for these cases. The estimated demand probability function $\Phi$ and its confidence bounds can be used to characterize the capability of a VB. For example, for the VB of this study, we can state that it can supply a power $x\in[-1138.3,5841.8]$ during $15$ minutes with probability greater than $0.98$ and confidence $0.995$. Moreover, we can also state from the data represented in Figure 3 that the VB can supply a power $x=5900$ kW during $15$ minutes with probability $p\in[0.7021,0.8322]$ and confidence $0.995$. Thus, the knowledge of the power supply probability function $\Phi$ for a given VB is crucial to decide about participating in electrical markets.
Table 2: Range of values for the parameters of residential TCLs [32] (pump set-points from [33]).
Heating pumps $2$ $2.0$ $-5.6$ $3.5$ $22.0$ $0.5$
Cooling pumps $2$ $2.0$ $5.6$ $2.5$ $24.0$ $0.5$
Electric water heater $120$ $0.4$ $-4.5$ $1.0$ $48.5$ $3.0$
Refrigerator $90$ $0.6$ $0.3$ $2.0$ $2.5$ $1.5$
4.2 Case study 2: Demand flexibility prediction
Table 3: Flexibility power available at each scenario (kW).
Scenario 1 Scenario 2 Scenario 3
3608.0 -1482.9 -1250.1
The resulting flexibility power at each scenario is showed in Table 3. The highest amount of power flexibility is obtained in Scenario 1, Followed by Scenario 2 and Scenario 3. Usually, the VB can provide more positive than negative power flexibility because TCLs take longer to increase or decrease their temperature when this change is not caused by an electromechanical element [9]. All calculated powers are above 1 MW in absolute value, which allows the aggregator to make bids in the SEBMs. These results demonstrate that the MC&ESB method is useful to calculate the availability of flexible power to participate in SEBMs and other similar markets in a short period of time, depending on the computational capabilities.
4.3 Case study 3: Demand flexibility control
Figure 5: Absolute error of the VB controller.
5 Conclusion
@@ A Bernouilli trial
Data availability
Data will be made available on request.
Conflicts of interest
The authors declare that they have no conflicts of interest.
Funding statement